Circular No. 5964

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994I IN NGC 5194 A. V. Filippenko, T. Matheson, and A. J. Barth, University of Cal-

ifornia at Berkeley, report: “CCD spectra (range 310–990 nm) obtained on Apr. 4.4 UT with the Shane 3-m reflector at Lick Observatory suggest that SN 1994I might be of type Ib rather than type II (cf. IAUC 5962), although the possible presence of weak Hα would technically make it ‘type IIb’ (cf. SN 1987K; Filippenko 1988, A.J. 96, 1941). The continuum drops quite steeply towards blue and near-ultraviolet wavelengths, unlike the case in most young type-II supernovae, but similar to the continua of type-Ib supernovae. Other aspects of the spectrum also support the type-Ib classification; in particular, lines of He I seem to be unusually strong, although this is not necessarily inconsistent with a young type-II object. Strong, narrow Na I D and Ca II H+K absorption lines suggest that the object is substantially reddened.”

G. C. L. Aikman and D. D. Balam, Dominion Astrophysical Observatory, report photometry obtained with the 1.82-m Plaskett telescope: Apr. 3.18 UT, V = 13.74 ± 0.06, B − V = +0.463, V − R = +0.134 (based on photometry by H. Corwin for the star given as ‘135’ on Thompson and Bryan’s Supernova Search Chart for NGC 5194: V = 13.44, B−V = +0.61, U − B = +0.09, V − R_{c} = +0.36). The effects of galaxy contamination were minimized by isophotal fitting and model galaxy subtraction.

Visual magnitude estimates: Apr. 1.25 UT, [14 (J. Leonard, Anza, CA); 2.27, – (P. Schwaar, Phoenix, AZ; independent visual discovery); 2.62, 14.2 (Y. Fujita, Kuma-Kogen Observatory, independent CCD discovery).

COMET MUELLER (1993a)

Ephemeris from orbital elements on MPC 23105:

1994TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Apr. 8 22^{h}54.^{m}67 − 1^{o}02.^{′}0 2.959 2.199 33.^{o}9 14.^{o}7 10.3
18 23 00.47 − 2 49.8 2.891 2.258  42.7 17.6 10.3
28 23 05.20 − 4 50.9 2.800 2.321  52.1 20.0 10.4
May 8 23 08.64 − 7 10.9 2.692 2.388  61.9 21.9 10.4
18 23 10.48 − 9 56.1 2.570 2.458  72.2 23.1 10.5
28 23 10.29 −13 13.0 2.443 2.531  83.2 23.4 10.5
June 7 23 07.56 −17 07.4 2.317 2.606  94.8 22.8 10.5
17 23 01.59 −21 43.0 2.204 2.684 107.0 21.2 10.5

1994 April 4 Daniel W. E. Green